OpenStudy (anonymous):

Help please! Thank you! (:

OpenStudy (anonymous):

Rationalize the denominator of sqrt -4/(7 - 3i)+(2 + 5i)

OpenStudy (anonymous):

\[\frac{\sqrt{-4}}{(7+3i)+(2+5i)}\]??

Parth (parthkohli):

Rationalize or "realize"?

Parth (parthkohli):

And first add the denominators, then conjugation...

OpenStudy (anonymous):

start with \[\frac{2i}{9+2i}\]

OpenStudy (anonymous):

Yes that is the equation

Parth (parthkohli):

I mean simplify.

Parth (parthkohli):

Right there, and then find the conjugate

OpenStudy (anonymous):

hold on one moment please

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks Parth

OpenStudy (anonymous):

And satelitte

Parth (parthkohli):

Yay. What was your answer?

OpenStudy (anonymous):

I got -16 + 18i/7

Parth (parthkohli):

Seems incorrect...

OpenStudy (anonymous):

ok one second lemme try again

OpenStudy (anonymous):

4 + 18i over 85

Parth (parthkohli):

That's right

OpenStudy (anonymous):

Thank you again!

Parth (parthkohli):

No problem at all. :-)